OpenStudy (anonymous):

(4 pt) What is the solution set for the given inequality if the replacement set for r is {5, 6, 7, 8, 9, 10}? 5r ≤ 6r – 8 A. {8, 9, 10} B. {5, 6, 7, 8} C. {9, 10} D. {5, 6, 7}

OpenStudy (anonymous):

@Jonathan1997

OpenStudy (anonymous):

B if r=8 2(8)≤ 3(8)-8 16≤ 16 if r=9 2(9)≤3(9)-8 18≤ 19 if r=10 2(10)≤ 3(10)-8

OpenStudy (anonymous):

a or d

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Which value is a solution to the open sentence? 4x – 5 = 31 A. 14 B. 6 C. 9 D. 12

OpenStudy (anonymous):

The answer is C